OpenStudy (anonymous):

Multiplaying by 3/4 and then dividing by 3/5 is the same thing as dividing by what?

OpenStudy (anonymous):

\[ \frac 34 \frac 5 3 =? \]

OpenStudy (anonymous):

\[ \frac 34 \frac 5 3 = \frac 5 4 \] It is like multiplying by \[ \frac 5 4 \] What would be the answer to your problem?

OpenStudy (anonymous):

4/5 is the correct answer, but I don't know how to get that.

OpenStudy (anonymous):

Multiplying by \[ \frac 5 4 \] is like dividing by \[ \frac 4 5 \]

OpenStudy (anonymous):

Oh! thanks so much!

OpenStudy (anonymous):

YW